# latexfontFont shape

### v2

@@ 4 Time-reversal symmetry and its breaking
Before dwelling into the construction of fluctuation theorems in the framework of general relativistic stochastic thermodynamics, it crucial to accurately describe what the TRS is meant on curved spacetime.
@@
To maintain general covariance in our construction, it is better to treat the TRT as a change of the observer’s time orientation, i.e. from the perspective of Alice to that of Carol, rather than treating it as a coordinate transformation. Meanwhile, we adopt the scalar expression like $-Z_{\mu}p^{\mu}$ for defining the energy of a particle instead of using the coordinate dependent definition $p^{0}$. Therefore, we employ an alternative representation of the TRT, which reverses the sign of full momentum (i.e. $p^{\mu}\mapsto-p^{\mu}$) while keeping the spacetime coordinates intact. Such a representation implies the reversal of the particle’s proper time derivative: ${\rm d}/{\rm d}\tau\mapsto-{\rm d}/{\rm d}\tau$. If Alice observes a particle evolving from $\tau_{I}$ to $\tau_{F}$ in its own proper time during the time interval $t_{I}\rightarrow t_{F}$, then Carol will observe the proper time of the charged particle evolves from $\tau_{F}$ to $\tau_{I}$.
Since the momentum is defined only in the local tangent space on curved spacetime, the TRT can be realized as an automorphism of $T_{x}\mathcal{M}$:
@@
Therefore, the phase trajectories lifted by Alice and Carol are on two disconnected regions of the mass shell bundle. For convenience, the trajectory lifted by Alice will be referred to as the forward trajectory, and the one by Carol will be referred to as the reversed trajectory. For later reference, we also list the induced action of the map $I$ on the coordinate basis for the vector field on $T\mathcal{M}$,
@@
where $\Sigma_{t}^{-}$ is the space of microstates associated to the past directed observer, Carol.
It remains to describe the TRT for the electromagnetic field. Consider the motion of a massive charged particle subjecting to the electromagnetic field $F^{\mu\nu}$. The equation of motion reads
@@
The TRS implies that the above equation of motion is kept invariant. Since $p^{\mu}$ is odd and hence $\frac{{\rm d}p^{\mu}}{{\rm d}\tau}$ is even under the TRT, the invariance of the equation of motion requires either
@@
Bringing all the above conventions together, we will see that the equations for the phase trajectory of a massive charged particle moving in a generic curved spacetime and subjecting to an external electromagnetic field is kept invariant under the TRT,
@@
Let us now bring the above system into a broader picture. Consider a scenario in which there are a great number of massive charged particles of different masses and charges moving together. The electromagnetic field is produced by the particles themselves and the spacetime geometry is determined by the masses and charges of the particles. We assume that the system is consisted of different species of particles, each species carries different mass $m_{s}$ and charge $q_{s}$ and obey a different distribution $\Phi_{s}$, which are differentiated from each other by the suffix $s$. We also assume that the distributions $\Phi_{s}$ are invariant under the TRT, i.e. $I^{*}\Phi_{s}=\Phi_{s}$. Then the total electric current
@@
should be both TRT invariants. Consequently, the Maxwell equation
@@
Stochastic mechanics can be regarded as an effective theory for describing a complete mechanical system within specific spatial and temporal scales [31, 32, 33]. Let us consider a heavy particle in the above system, assuming that the remaining particles constitute a heat reservoir which has already reached equilibrium. The electromagnetic interaction acting on the heavy particle can be divided into two parts: the coarse-grained averaging effects at larger spatial and temporal scales, and the stochastic remnants at smaller scales. Consequently, eqs. (75)-(76) can be approximated by LEτ [1], and this heavy particle is the Brownian particle.
However, such a coarse-grained description violates the TRS. Let us first consider the damping force. If the forward and reversed trajectories simultaneously satisfy following equation
@@
where $Q_{\rm B}[\varphi]$ is the heat transfer rate from Bob’s perspective, and $Q_{R}$ is that of the heat reservoir. It is worth noticing that the Clausius’ identity holds only from the perspective of Bob.
@@ 5 Fluctuation theorem
where $\tilde{Y}_{[t]}=(\tilde{Y}_{0},\tilde{Y}_{1},\cdots,\tilde{Y}_{N})$ and $\tilde{Y}_{[t]}^{-}=(\tilde{Y}^{-}_{0},\tilde{Y}_{1}^{-},\cdots,\tilde{Y}_{N}^{-})$ denote the forward and reversed processes, and $Y_{[t]}$ and $Y^{-}_{[t]}$ denote the forward and reversed trajectories which are already described in Sec. 4. Notice that “forward process” and “forward trajectory” are different concepts: the latter is a concrete realization of the former. Unlike the relation between forward and reversed trajectories, the only requirement on the reversed process is that its initial state is identical to the TRT of the final state of the forward process, i.e.
@@
In order to describe the fluctuation theorems at the trajectory level, we need a discretized version of LEt, which reads
@@
where $\tilde{Y}_{\bar{n}}:=(\tilde{Y}_{n+1}+\tilde{Y}_{n})/2$ which comes from the Stratonovich coupling. We also need to introduce the measures on the space of forward and reversed trajectories which is defined to be the successive wedge products of the volume element of the space of microstates at each moment of the discrete time,
@@
$=\Delta S_{R}+\Delta S.$ (111)
Please be reminded that the relation eq. (30) between different distribution functions and the definition (60) of the trajectory entropy production are used here. Denoting the total entropy production at the trajectory level as $\Sigma_{Y_{[t]}}:=\Delta S_{R}+\Delta S$, the usual form of the detailed fluctuation theorem arises
This theorem reveals how the breaking of the TRS is associated with the entropy production at the trajectory level. Moreover, taking the expectation value of ${\rm e}^{\Sigma_{X_{[t]}}}$ and using Jensen inequality, one gets the integral fluctuation theorem
${\rm e}^{-\left\langle\Sigma_{\tilde{Y}_{[t]}}\right\rangle}\leq\left\langle{\rm e}^{-\Sigma_{\tilde{Y}_{[t]}}}\right\rangle=\int\mathcal{D}[Y_{[t]}]\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]=\int\mathcal{D}[Y^{-}_{[t]}]\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]=1,$ (113)
which implies that the statistical expectation value of the total entropy production must be no-negative, i.e. $\left\langle\Sigma_{\tilde{Y}_{[t]}}\right\rangle\geq 0$.
@@ 6 Concluding remarks
Fluctuation theorems not only are important in understanding the theoretical origin of irreversibility, but also have found various applications in diverse areas ranging from macroscopic to mesoscopic systems, and even to active matter. The research carried our in the present work opens new area for the potential applications of fluctuation theorems which embodies relativistic gravity, as in typical cases of cosmological and astrophysical processes.
In our construction, it is important to decouple the choice of observers from the choice of coordinate systems, which has several important consequences. First, it allows us to parametrize the stochastic trajectory using the proper time of the observer field, which helps to get rid of the random clock carried by the Brownian particle. Second, it helps to properly describe the TRT in a coordinate independent manner. Last, it is precisely such decoupling that makes the construction fully general covariant.
Although the values of many thermodynamic quantities are observer dependent, including e.g. the energy and its density, the temperature and chemical potential, and even the Clausius’ identity etc., the total entropy production is not among the observer dependent quantities. Meanwhile, the trajectory probability is a pure mathematical entity which is also observer independent. Therefore, it is not surprising that the form of the fluctuation theorems obtained in this work is identical to that obtained from non-relativistic stochastic thermodynamics.
@@ Appendix A Coordinate bases for tangent vectors
In our description of the tangent bundle $T\mathcal{M}$, we employ $22d+$ 2 independent coordinates $(x^{\mu},p^{\mu})$ with the corresponding coordinate basis $\left(\left.\frac{\partial}{\partial x^{\mu}}\right|_{T\mathcal{M}},\left.\frac{\partial}{\partial p^{\mu}}\right|_{T\mathcal{M}}\right)$ for tangent vectors on $T\mathcal{M}$. However, due to the mass shell constraint, the above coordinates are no longer independent on the mass shell bundle $\Gamma^{+}_{m}$, rendering the coordinate basis also redundant. The same also happens when we restrict ourselves from the mass shell bundle to the space of microstates $\Sigma_{t}^{+}$. The aim of this appendix is to resolve the redundancies in the coordinate bases on various useful submanifolds of $T\mathcal{M}$.
@@
Therefore, the tangent vectors of the momentum space have two different representations under the above two bases. We will use them interchangeably. There is a similar property for the vectors on mass shell bundle. Let $\mathscr{V}$ be a tangent vector of the phase trajectory $X_{\tau}:=(x_{\tau},p_{\tau})$. As a vector on tangent bundle, $\mathscr{V}$ can be expanded as
@@
The space of microstates $\Sigma^{+}_{t}$ is also an embedding submanifold of the mass shell bundle defined by fixing $t(x)$ to be a constant. Therefore, the coordinates on $\Gamma^{+}_{m}$ becomes redundant once again on $\Sigma^{+}_{t}$. Such redundancy can also be eliminated by taking $(x^{i},p^{i})$ as coordinate of $\Sigma^{+}_{t}$ and viewing $x^{0}(t,x^{i})$ as a function. The restriction of $f$ on $\Sigma^{+}_{t}$ is defined as
@@ Appendix B Continuum limit
1. Let $A$ be a full rank square matrix and $B$ be an arbitrary matrix of the same size. Then the determinant of $A+B{\rm d}t+o(t^{2})$ can be expanded into power series in ${\rm d}t$,
@@
2. Let $f(t)$ be a continuous function on $[t_{I},t_{F}]$, then
@@
Let $\mathscr{O}$ be a function on the mass shell bundle and $Y_{t}=X_{\tau(t)}$ be a phase trajectory parametrized by the proper time of Alice. Then
@@
The generalization of this equation to general dimension is straightforward, telling that the probability distribution of $\tilde{y}^{i}_{n+1}$ under the conditions $\tilde{k}^{i}_{n+1}={k}^{i}_{n+1},\tilde{Y}_{n}=Y_{n}$ is given by
@@
The appearance of $\partial/\partial\hat{y}^{i}$ in the second line of eq. (B) is due to the fact that $\tilde{y}^{i}_{n+1}$ only takes values in the configuration space, and its probability density is a scalar field on $\mathcal{S}_{n}$. The factor $|\partial_{0}t/(\lambda\sqrt{g})|$ arises from the change of coordinate volume element into invariant volume element, as indicated by eq. (2).
@@
where $\mathcal{A}^{i}_{n}:=\gamma|_{Y_{\bar{n}}}(k^{i}_{n+1}-k^{i}_{n})/(m{\rm d}t)$ is the coordinate acceleration, then eq. (102) can be rewritten as
@@
In the case that the realization of $\tilde{Y}_{n}$ is given, $\tilde{y}_{n+1}$ is determined by $\tilde{k}_{n+1}$. Therefore, eq. (147) gives the relation between ${\rm d}\tilde{W}^{\mathfrak{a}}_{n}$ and $\tilde{k}^{i}_{n+1}$, which can be used to calculate the conditional probability
@@
which is exactly eq. (5) appeared in the main text.

### v3

@@ 4 Time-reversal symmetry and its breaking
Before delving into the construction of fluctuation theorems in the framework of general relativistic stochastic thermodynamics, it is crucial to accurately describe what the TRS is meant on curved spacetime.
@@
To maintain general covariance in our construction, it is better to treat the TRT as a change of the observer’s time orientation, i.e. from the perspective of Alice to that of Carol, rather than treating it as a coordinate transformation. Meanwhile, we adopt the scalar expression like $-Z_{\mu}p^{\mu}$ for defining the energy of a particle instead of using the coordinate dependent definition $p^{0}$. Therefore, we employ an alternative representation of the TRT, which reverses the sign of full momentum (i.e. $p^{\mu}\mapsto-p^{\mu}$) while keeping the spacetime coordinates intact. Such a representation implies the reversal of the particle’s proper time derivative: ${\rm d}/{\rm d}\tau\mapsto-{\rm d}/{\rm d}\tau$. If Alice observes a particle evolving from $\tau_{I}$ to $\tau_{F}$ in its own proper time during the time interval $t_{I}\rightarrow t_{F}$, then Carol will observe the proper time of the charged particle evolving from $\tau_{F}$ to $\tau_{I}$.
Since the momentum is defined only in the local tangent space on curved spacetime, the TRT can be realized as an automorphism of $T_{x}\mathcal{M}$ (see Fig.2):
@@
Therefore, the phase trajectories lifted by Alice and Carol are on two disconnected regions of the mass shell bundle. For convenience, the trajectory lifted by Alice will be referred to as the forward trajectory, and the one lifted by Carol will be referred to as the reversed trajectory. For later reference, we also list the induced action (i.e. pushforward) of the map $I$ on the coordinate basis for the vector field on $T\mathcal{M}$,
@@
where $\Sigma_{t}^{-}$ is the space of microstates associated with the past directed observer, Carol.
It remains to describe the TRT for the electromagnetic field. Consider the motion of a massive charged particle subjected to the electromagnetic field $F^{\mu\nu}$. The equation of motion reads
@@
The TRS implies that the above equation of motion remains invariant. Since $p^{\mu}$ is odd and hence $\frac{{\rm d}p^{\mu}}{{\rm d}\tau}$ is even under the TRT, the invariance of the equation of motion requires either
@@
Bringing all the above conventions together, we will see that the equations for the phase trajectory of a massive charged particle moving in a generic curved spacetime and subjected to an external electromagnetic field remains invariant under the TRT,
@@
Let us now bring the above system into a broader picture. Consider a scenario in which there are a great number of massive charged particles of different masses and charges moving together. The electromagnetic field is produced by the particles themselves and the spacetime geometry is determined by the masses and charges of the particles. We assume that the system consists of different species of particles, and each species carries different mass $m_{s}$ and charge $q_{s}$ and obeys a different distribution $\Phi_{s}$, which are differentiated from each other by the suffix $s$. We also assume that the distributions $\Phi_{s}$ are invariant under the TRT, i.e. $I^{*}\Phi_{s}=\Phi_{s}$. Then the total electric current
@@
should both be TRT invariants. Consequently, the Maxwell equation
@@
Stochastic mechanics can be regarded as an effective theory for describing a complete mechanical system within specific spatial and temporal scales [32, 33, 34]. Let us consider a heavy particle in the above system, assuming that the remaining particles constitute a heat reservoir that has already reached equilibrium. The electromagnetic interaction acting on the heavy particle can be divided into two parts: the coarse-grained averaging effects at larger spatial and temporal scales, and the stochastic remnants at smaller scales. Consequently, eqs. (75)-(76) can be approximated by LEτ [19], and this heavy particle is the Brownian particle.
However, such a coarse-grained description violates the TRS. Let us first consider the damping force. If the forward and reversed trajectories simultaneously satisfy the following equation
@@
where $Q_{\rm B}[\varphi]$ is the heat transfer rate from Bob’s perspective, and $Q_{R}$ is that of the heat reservoir. It is worth noticing that Clausius’ identity holds only from the perspective of Bob.
@@ 5 Fluctuation theorem
where $\tilde{Y}_{[t]}=(\tilde{Y}_{0},\tilde{Y}_{1},\cdots,\tilde{Y}_{N})$ and $\tilde{Y}_{[t]}^{-}=(\tilde{Y}^{-}_{0},\tilde{Y}_{1}^{-},\cdots,\tilde{Y}_{N}^{-})$ denote the forward and reversed processes, and $Y_{[t]}$ and $Y^{-}_{[t]}$ denote the forward and reversed trajectories which are already described in Sec. 4. Notice that “forward process” and “forward trajectory” are different concepts: the latter is a concrete realization of the former. Unlike the relation between forward and reversed trajectories, the only requirement in the reversed process is that its initial state is identical to the TRT of the final state of the forward process, i.e.
@@
Using the above conventions, we are now in the right position to present the precise form of the detailed and integral fluctuation theorems in the context of general relativistic stochastic thermodynamics and formulate their proofs.
Theorem 5.1 (Detailed fluctuation theorem)
The ratio between the probabilities for the forward and reversed general relativistic stochastic trajectories $Y_{[t]}$ and $Y^{-}_{[t]}$ to be realized is equal to the exponential of total entropy production $\Sigma_{Y_{[t]}}$ along the trajectory $Y_{[t]}$, i.e.
Theorem 5.2 (Integral fluctuation theorem)
The statistical expectation value of the total trajectory entropy production must be non-negative, i.e.
${\rm e}^{-\left\langle\Sigma_{Y_{[t]}}\right\rangle}\leq 1,\qquad\left\langle\Sigma_{Y_{[t]}}\right\rangle\geq 0.$ (96)
The rest of this section is devoted to the proof of the above theorems.
@@
In order to prove the fluctuation theorems at the trajectory level, we need a discretized version of LEt, which reads
@@
where $\tilde{Y}_{\bar{n}}:=(\tilde{Y}_{n+1}+\tilde{Y}_{n})/2$ which comes from the Stratonovich coupling. We also need to introduce the measures on the space of forward and reversed trajectories, which are defined to be the successive wedge products of the volume element of the space of microstates at each moment of the discrete time,
@@
$=\Delta S_{R}+\Delta S{\color[rgb]{0,0,0}~{}:=\Sigma_{Y_{[t]}}}.$ (113)
It is important to remember that the relation eq. (30) between different distribution functions and the definition (60) of the trajectory entropy production are used here. Exponentiating the above relation yields Theorem 5.1. Moreover, Theorem 5.2 follows immediately by use of Theorem 5.1 and taking an integration in the space of trajectories, together with the aid of Jensen inequality,
${\rm e}^{-\left\langle\Sigma_{Y_{[t]}}\right\rangle}\leq\left\langle{\rm e}^{-\Sigma_{Y_{[t]}}}\right\rangle$ $:=\int\mathcal{D}[Y_{[t]}]\Pr[\tilde{Y}_{[t]}=Y_{[t]}]~{}{\rm e}^{-\Sigma_{Y_{[t]}}}$
$=\int\mathcal{D}[Y^{-}_{[t]}]\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]=1,$ (114)
which implies $\left\langle\Sigma_{Y_{[t]}}\right\rangle\geq 0$.
@@ 6 Concluding remarks
We have thus proved the detailed and integral fluctuation theorems in the context of general relativistic stochastic thermodynamics. In our construction, it is important to decouple the choice of observers from the choice of coordinate systems, which has several important consequences. First, it allows us to parametrize the stochastic trajectory using the proper time of the observer field, which helps to get rid of the random clock carried by the Brownian particle. Second, it helps to properly describe the TRT in a coordinate independent manner. Lastly, it is precisely such decoupling that makes the construction fully general covariant.
Although the values of many thermodynamic quantities are observer dependent, including e.g. the energy and its density, the temperature and chemical potential, and even the Clausius’ identity etc., the total entropy production is not among the observer dependent quantities. Meanwhile, the trajectory probability is a pure mathematical entity that is also observer independent. Therefore, it is not surprising that the form of the fluctuation theorems obtained in this work is identical to that obtained from non-relativistic stochastic thermodynamics.
Fluctuation theorems are not only important in understanding the theoretical origin of irreversibility, but they have also found various applications in diverse areas ranging from macroscopic to mesoscopic systems, and even to active matter. For example, there are some studies [35, 36, 37, 38] utilizing the non-relativistic fluctuation theorem to explore the concept of Maxwell’s demon. Some of them argue that the change of the mutual information between Maxwell’s demon and the Brownian particle should be considered as a part of the entropy production. When considering the event horizon within curved spacetime, which serves as a causal boundary, the causal connection between Maxwell’s demon and the Brownian particle may be lost over time. Consequently, behaviors different from those predicted by non-relativistic fluctuation theorems, similar to the black hole information paradox, may emerge. Another possible scenario in which our general relativistic fluctuation theorems may find application is in black hole physics. In recent years, due to the observational progresses brought about by the event horizon telescope [39], a considerable amount of theoretical works on the images and shadows of black holes have appeared [^1] A quick search of the literature [43] has found 3222 results. It is impossible to list them all, therefore we simply refer to [40, 41, 42] as prototypes of related works.. In most of those works, the light-emitting substance is assumed to be particles undergoing regular geodesic motion, rendering the corresponding images also regular and containing rich patterns. In reality, however, since a large amount of particles accrete around the black holes, they effectively form a heat reservoir with a very high temperature, and consequently they make the trajectories of individual particles moving inside the accretion disk probabilistic. The corresponding images should also be obscured as a consequence of such probabilistic motion. The quantitative nature of the detailed fluctuation theorem may help to understand to what extent the black hole images should be obscured. In summary, the research carried out in the present work opens a new area for the potential applications of fluctuation theorems which embodies relativistic gravity, as in typical cases of cosmological processes, astrophysical processes and black hole physics. We believe that much more sophisticated applications of the new version of the fluctuation theorems presented in this work and more generally of the framework of general relativistic stochastic thermodynamics will emerge as we look more clearly and deeply into them. We are working hard on related problems.
@@ Appendix A Coordinate bases for tangent vectors
In our description of the tangent bundle $T\mathcal{M}$, we employ $22d+$ 2 independent coordinates $(x^{\mu},p^{\mu})$ with the corresponding coordinate basis $\left(\left.\frac{\partial}{\partial x^{\mu}}\right|_{T\mathcal{M}},\left.\frac{\partial}{\partial p^{\mu}}\right|_{T\mathcal{M}}\right)$ for tangent vectors on $T\mathcal{M}$. However, due to the mass shell constraint, the above coordinates are no longer independent of the mass shell bundle $\Gamma^{+}_{m}$, rendering the coordinate basis also redundant. The same also happens when we restrict ourselves from the mass shell bundle to the space of microstates $\Sigma_{t}^{+}$. The aim of this appendix is to resolve the redundancies in the coordinate bases on various useful submanifolds of $T\mathcal{M}$.
@@
Therefore, the tangent vectors of the momentum space have two different representations under the above two bases. We will use them interchangeably. There is a similar property for the vectors on the mass shell bundle. Let $\mathscr{V}$ be a tangent vector of the phase trajectory $X_{\tau}:=(x_{\tau},p_{\tau})$. As a vector on a tangent bundle, $\mathscr{V}$ can be expanded as
@@
The space of microstates $\Sigma^{+}_{t}$ is also an embedding submanifold of the mass shell bundle defined by fixing $t(x)$ to be a constant. Therefore, the coordinates on $\Gamma^{+}_{m}$ become redundant once again on $\Sigma^{+}_{t}$. Such redundancy can also be eliminated by taking $(x^{i},p^{i})$ as a coordinate of $\Sigma^{+}_{t}$ and viewing $x^{0}(t,x^{i})$ as a function. The restriction of $f$ on $\Sigma^{+}_{t}$ is defined as
@@ Appendix B Continuum limit
1. Let $A$ be a full rank square matrix, and let $B$ be an arbitrary matrix of the same size. Then the determinant of $A+B{\rm d}t+o(t^{2})$ can be expanded into power series in ${\rm d}t$,
@@
2. Let $f(t)$ be a continuous function on $[t_{I},t_{F}]$, and then
@@
Let $\mathscr{O}$ be a function on the mass shell bundle, and let $Y_{t}=X_{\tau(t)}$ be a phase trajectory parametrized by the proper time of Alice. Then
@@
The generalization of this equation to general dimension is straightforward, showing that the probability distribution of $\tilde{y}^{i}_{n+1}$ under the conditions $\tilde{k}^{i}_{n+1}={k}^{i}_{n+1},\tilde{Y}_{n}=Y_{n}$ is given by
@@
The appearance of $\partial/\partial\hat{y}^{i}$ in the second line of eq. (B) is due to the fact that $\tilde{y}^{i}_{n+1}$ only takes values in the configuration space, and its probability density is a scalar field on $\mathcal{S}_{n}$. The factor $|\partial_{0}t/(\lambda\sqrt{g})|$ arises from the change of coordinate volume element into an invariant volume element, as indicated by eq. (2).
@@
where $\mathcal{A}^{i}_{n}:=\gamma|_{Y_{\bar{n}}}(k^{i}_{n+1}-k^{i}_{n})/(m{\rm d}t)$ is the coordinate acceleration, eq. (104) can be rewritten as
@@
In the case in which the realization of $\tilde{Y}_{n}$ is given, $\tilde{y}_{n+1}$ is determined by $\tilde{k}_{n+1}$. Therefore, eq. (148) gives the relation between ${\rm d}\tilde{W}^{\mathfrak{a}}_{n}$ and $\tilde{k}^{i}_{n+1}$, which can be used to calculate the conditional probability
@@
which is exactly eq. (5) in the main text.